# Wilson Line and Disorder Invariants of Topological One-Dimensional Multiband Models

### v1

@@ III.3 SSH4 Model
$\mathcal{H}_{\text{SSH{\hbox{_{4}}}}}=\begin{pmatrix}0&t_{1}&0&t_{4}e^{-ik}\\ t_{1}&0&t_{2}&0\\ 0&t_{2}&0&t_{3}\\ t_{4}e^{ik}&0&t_{3}&0\end{pmatrix},$ (26)
@@ III.3.2 Wilson Line Invariant $\gamma$
$\gamma=\begin{cases}3&\left|\frac{t_{1}t_{3}}{t_{2}t_{4}}\right|<1\text{ and }\left|\frac{t_{2}t_{3}}{t_{1}t_{4}}\right|<1\\ 2&\left|\frac{t_{1}t_{3}}{t_{2}t_{4}}\right|>1\text{ and }\left|\frac{t_{2}t_{3}}{t_{1}t_{4}}\right|<1\\ 1&\left|\frac{t_{1}t_{3}}{t_{2}t_{4}}\right|<1\text{ and }\left|\frac{t_{2}t_{3}}{t_{1}t_{4}}\right|>1\\ 0&\left|\frac{t_{1}t_{3}}{t_{2}t_{4}}\right|>1\text{ and }\left|\frac{t_{2}t_{3}}{t_{1}t_{4}}\right|>1\end{cases}$ (29)
In this case, $\gamma$ can take integer values from 0 to 3, which predicts more edge states than indicated by either the polarization or the winding number. Fig. 5 illustrates an example where the unwrapped Wilson line yields an invariant of 3, in contrast to the polarization, which gives a value of 1.
@@
Numerically, the left and right edge states in Figs. 6a and 6c are computed by defining an edge state when the localization length of the edge state is smaller than the system length ($L_{c}<N$), where $L_{c}^{-1}$ is the average slope of $\ln(|\psi_{i}|)$ along the finite chain indexed by site $i$. This slope is negative for a left-edge state and positive for a right-edge state. Since the edge states are inside the gap, it is also possible to extract them by looking at the DOS. This method counts the sum of left and right edge states. The Wilson line invariant corresponds to the topology obtained from edge states in the limit of $N\rightarrow\infty$.
@@ III.3.3 Robustness of Topological Edge States to Disorder
To further identify the topological nature of the edge states, we can look at their behavior as a function of disorder. The $E=0$ edge states are protected by chiral symmetry preserving disorder, such as off-diagonal disorder ($t_{i}\rightarrow t_{i}+V_{\text{dis}}$). We characterize the strength of disorder with $W$, where the random variable, $V_{\text{dis}}$, is uniformly distributed within $-W<V_{\text{dis}}<W$. In contrast to the $E=0$ edge states, the non-zero energy edge states ($E_{\pm}$) are not protected by conventional chiral symmetry but rather by a hidden chiral symmetry, dubbed $C_{1/2}$ [35]. For $C_{1/2}$ symmetry to protect the in-gap states, the SSH4 chain must have the four hopping amplitudes arranged so that the following conditions are met. First that the chain remains bipartite and secondly no cyclic shift of the hoppings makes the sequence palindromic ($t_{1}\neq t_{3}$ and $t_{2}\neq t_{4}$), so that the physical inversion axis cannot be centered in any unit cell.
@@
Figure 8: wavefunction amplitude ($\psi$) distributed across the lattice sites, where we only show the 16 left-most and right-most sites of a total of 2000 sites. In the top graph we show $E=0$ edge sate solutions for 100 different off-diagonal disorder configurations with a disorder strength of $W=0.25$. Here we chose a general off-diagonal disorder distributed over all the hopping elements. In the bottom graph, we used onsite D disorder with strength $W=1$ for 100 different disorder configurations for the $E=E_{+}$ solutions. The 100 left Edge solutions are identical (only one line is visible), unlike the 100 right edge solutions (multiple lines). We used the same parameters as in figure 7 with $t_{2}=1.105$ for the top graph and $t_{2}=-0.05$ for the bottom graph.
@@ IV Conclusion
We have studied the topological invariant for the EnSSH and SSH4 models in some detail. However, our results can be generalized to other models such as SSHn, where $n$ is even. In this case, the number of bands is given by $n$ and there can be $n$-1 band gaps, with corresponding edge states. It is also possible to add long range hopping to the SSHn model, which we label EmSSHn, where $m$ and $n$ can be arbitrary integers. While the general solutions can quickly become quite involved for large $m$ and $n$, calculating the $\mathbb{Z}$ topological invariant using the Wilson line approach we introduced in equ. (17) is numerically straightforward and allows one to label all the topological phases with their corresponding invariants. We should also note that the 1D models considered here are somewhat special in the sense that they are often described as boundary obstructed topological phases [38] in contrast to topological phases indexed by Chern numbers. Non-zero Chern numbers do not exist in 1D.
Extensions to higher dimensions are also possible in the context of SSH-type models [6], which also lead to rich topological phases [8]. Here, we restricted ourselves to extended 1D models, which also have a similarly rich topological phase diagram with multiple invariants (see for example, figs. 3 and 4). We provide here a clear connection between the topological invariant and the finite-size edge-state picture using the bulk-boundary correspondence. A finite-sized system, as relevant for experimental systems, allows the identification of the topological invariant, but not the sign change. To circumvent this limitation, we propose a new configuration, which would also allow one to probe of the sign of the topological invariant in extended SSH models. We further make a clear connection between the symmetries of the edge states and their robustness to disorder. It is indeed possible to ”kill” some of the edge states with disorder while preserving other ones in the same system. This extends our toolbox even further to probe and manipulate topological states.

### v2

@@ III.3 SSH4 Model
$\mathcal{H}_{\text{SSH${}_{4}$}}=\begin{pmatrix}0&t_{1}&0&t_{4}e^{-ik}\\ t_{1}&0&t_{2}&0\\ 0&t_{2}&0&t_{3}\\ t_{4}e^{ik}&0&t_{3}&0\end{pmatrix}$, (26)
@@ III.3.2 Wilson Line Invariant $\gamma$
[…]
In this case, $\gamma$ can take integer values from 0 to 3, which predicts more edge states than indicated by polarization or winding number. Fig. 5 illustrates an example, where the unwrapped Wilson line yields an invariant of 3, in contrast to the polarization, which gives a value of 1.
@@
Numerically, the left and right edge states in Figs. 6a and 6c are computed by defining an edge state when the localization length of the edge state is smaller than the system length ($L_{c}<N$), where $L_{c}^{-1}$ is the average slope of $\ln(|\psi_{i}|)$ along the finite chain indexed by site $i$. This slope is negative for a left-edge state and positive for a right-edge state. Since the edge states are inside the gap, it is also possible to extract them by looking at the DOS. This method counts the sum of the left and right edge states. The Wilson line invariant corresponds to the topology obtained from edge states in the limit of $N\rightarrow\infty$.
@@ III.3.3 Robustness of Topological Edge States to Disorder
To further identify the topological nature of the edge states, we can look at their behavior as a function of disorder. The $E=0$ edge states are protected by chiral symmetry preserving disorder, such as off-diagonal disorder ($t_{i}\rightarrow t_{i}+V_{\text{dis}}$). We characterize the strength of the disorder with $W$, where the random variable, $V_{\text{dis}}$, is uniformly distributed within $W<V_{\text{dis}}<W$. In contrast to the $E=0$ edge states, the non-zero energy edge states ($E_{\pm}$) are not protected by conventional chiral symmetry but rather by a hidden chiral symmetry, dubbed $C_{1/2}$ [37]. For $C_{1/2}$ symmetry to protect the in-gap states, the SSH4 chain must have the four hopping amplitudes arranged so that the following conditions are met. First that the chain remains bipartite and secondly no cyclic shift of the hoppings makes the sequence palindromic ($t_{1}\neq t_{3}$ and $t_{2}\neq t_{4}$), so that the physical inversion axis cannot be centered in any unit cell.
@@
Figure 8: wavefunction amplitude ($\psi$) distributed across the lattice sites, where we only show the 16 left-most and right-most sites of a total of 2000 sites. In the top graph we show $E=0$ edge state solutions for 100 different off-diagonal disorder configurations with a disorder strength of $W=0.25$. Here we chose a general off-diagonal disorder distributed over all the hopping elements. In the bottom graph, we used onsite D disorder with strength $W=1$ for 100 different disorder configurations for the $E=E_{+}$ solutions. The 100 left Edge solutions are identical (only one line is visible), unlike the 100 right edge solutions (multiple lines). We used the same parameters as in figure 7 with $t_{2}=1.105$ for the top graph and $t_{2}=-0.05$ for the bottom graph.
@@ IV Conclusion
We have studied the topological invariant for the EnSSH and SSH4 models in some detail. Our results can be generalized to other models such as SSHn as long as each band is gapped. In this case, the number of bands is given by $n$ and nn-1 band gaps with corresponding edge states. The case of the SSH3 has been extensively studied in [12] and provides a good discussion on a gauge invariant topological invariant and protecting symmetries. It is also possible to add long range hopping to the SSHn model, which we label EmSSHn, where $m$ and $n$ can be arbitrary integers. While the general solutions can quickly become quite involved for large $m$ and $n$, calculating the $\mathbb{Z}$ topological invariant using the Wilson line approach we introduced in equ. (17) is numerically straightforward and allows one to label all the topological phases with their corresponding invariants. We should also note that the 1D models considered here are somewhat special in the sense that they are often described as boundary obstructed topological phases [40] in contrast to topological phases indexed by Chern numbers. Non-zero Chern numbers do not exist in 1D.
Extensions to higher dimensions are also possible in the context of SSH-type models [6], which also lead to rich topological phases [8]. Here, we restricted ourselves to extended 1D models, which also have a similarly rich topological phase diagram with multiple invariants (see for example, figs. 3 and 4). We provide here a clear connection between the topological invariant and the finite-size edge-state picture using the bulk-boundary correspondence. A finite-sized system, as relevant for many experimental systems, allows the identification of the topological invariant, but not the sign change. To circumvent this limitation, we propose a new configuration, which would also allow one to probe of the sign of the topological invariant in extended SSH models. We further make a clear connection between the symmetries of the edge states and their robustness to disorder. It is indeed possible to ”kill” some of the edge states with disorder while preserving other ones in the same system. This extends our toolbox even further to probe and manipulate topological states.